OpenStudy (anonymous):

2) WILL GIVE MEDAL!! Use the quadratic formula to solve 2y2 – 5y = –2. a. {–½, 2} b. {–½, –2} c. {½, 2} d. {½, –2} @wolf1728

OpenStudy (wolf1728):

a=2 b=-5 c=2 y = [--5 +-sqroot(25 -4*2*2)] / 4 y = [5 +sqroot(9)]/4 y1 = 2 y2 = [--5 -sqroot(25 -4*2*2)] / 4 y2 = (5 -3) / 4 y2 = .5

OpenStudy (wolf1728):

Do you see why c=2 and NOT c=-2?

OpenStudy (anonymous):

kinda its because it has to be a positive...

OpenStudy (anonymous):

so it would be A

OpenStudy (wolf1728):

No. In order for the quadratic equation to work, the equation must always be set up ay^2 +by +c = 0 and the answer is c

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

ohh ok!!